OpenStudy (anonymous):

A cylindrical can is 20cm in diameter and 16cm in height. You want to reduce the diameter of the can to 16cm. What must the height be if the new can still has the same volume?

OpenStudy (anonymous):

\[Volume (Cylinder) = \pi*r^2h\] Therefore, if we change the radius by x/y, we must change the height by y^2/x^2 Change in radius: 16/20 = 4/5 Change in height: 25/16 New height: 25cm

OpenStudy (anonymous):

Thank you for explaining! :) I really needed help!